snowflake0531:

not sure about 9 and 14 @AZ

AZ:

another way to write it would be \(\Large\sqrt[3]{x} = x^{\frac{1}{3}}\)

AZ:

\(\Large\log_a b^c = c ~\log_a b\)

snowflake0531:

so \[ \frac{1}{3} ln ~x\]

AZ:

see, how hard was that?

snowflake0531:

e.e i forgot about fraction exponents *facepalm

AZ:

the second one, you have to remember \(\Large \log (ab) = \log a + \log b\)

snowflake0531:

And then for 14 just 5log10 + log sqrty ?

AZ:

and then use

@az wrote:
\(\Large\log_a b^c = c ~\log_a b\)

AZ:

also log 10 = 1 and what can we write log sqrt(y) as? hint: think about what we did to #9

snowflake0531:

5 + 1/2 log y

AZ:

ta-da

snowflake0531:

tyty xdd

AZ:

ywyw xdd